OpenStudy (anonymous):

Sally buys 5 shirts on sale for $20 each. She also buys 5 pairs of shorts, each pair of shorts being the same price. If the total cost for the clothes is $185, how much did each pair of shorts cost?

OpenStudy (anonymous):

If you multiply 5 by 20 you get 100 then subtract that form 185 giving you 85 and divide by 5 to find the unit price of a pair of shorts which in this case is $17

OpenStudy (texaschic101):

s = shirts = 20 y = shorts =========== 5s +5y = 185 --- 5 shirts + 5 shorts = 185) 5(20) + 5y = 185 100 + 5y = 185 5y = 185 - 100 5y = 85 y = 85/5 y = 17 yep....each pair of shorts costs $ 17